OpenStudy (anonymous):

in pythagorean therom if you have a and c but no b you subtract not add right?

OpenStudy (lgbasallote):

\[a^2 + b^2 = c^2\] \[b^2 = c^2 - a^2\] \[b = \sqrt{c^2 - a^2}\] does that answer your question?